### v1

@@ 3 PageRank problem
Assume $d(v)\geq 1$ for all $v\in\mathcal{V}$. For a proper subset $\mathcal{C}$ of $\mathcal{V}$ with $s\in\mathcal{C}$,
$x[\bar{\mathcal{C}}]<\frac{1}{\beta}w\left(\partial(\mathcal{C})\right).$
Proof 3.6.
A summation of $x(v)$ over all $v\in\bar{\mathcal{C}}$ equates to
$x[\bar{\mathcal{C}}]=\sum_{v\in\bar{\mathcal{C}}}\frac{1}{\left(\beta d(v)+1\right)d(v)}\sum_{u\sim v}w(u,v)x(u).$
Since $d(v)\geq 1$ for all $v\in\mathcal{V}$, it follows that
$x[\bar{\mathcal{C}}]$ $\leq\frac{1}{\beta+1}\sum_{v\in\bar{\mathcal{C}}}\sum_{u\sim v}w(u,v)x(u)$
$=\frac{1}{\beta+1}\left(\sum_{\begin{subarray}{c}u\sim v\\ u,v\in\bar{\mathcal{C}}\end{subarray}}w(u,v)\left(x(u)+x(v)\right)+\sum_{\begin{subarray}{c}u\sim v\\ u\in\mathcal{C},v\in\bar{\mathcal{C}}\end{subarray}}w(u,v)x(u)\right).$
Substitution of the first summation term on the right-hand side with $x[\bar{\mathcal{C}}]$ yields
$x[\bar{\mathcal{C}}]\leq\frac{1}{\beta}\sum_{\begin{subarray}{c}u\sim v\\ u\in\mathcal{C},v\in\bar{\mathcal{C}}\end{subarray}}w(u,v)x(u).$
By virtue of Proposition 3.3 and $d(s)\geq 1$,
$x[\bar{\mathcal{C}}]<\frac{1}{\beta}w\left(\partial(\mathcal{C})\right).$
4 Nonlinear Modified PageRank problem
@@
is proposed for local graph partitioning purposes. $f$ is required to be continuous and have at least continuous first-order partial derivatives with respect to the entries of $x$. This begs the question as to what form should $f$ take.
@@
Let $p\in[1,\infty)$. If $\frac{\partial}{\partial x}\left(\lVert Bx\rVert_{p}^{p}\right)=0$, then
(8) $B^{\top}\left(|Bx|^{\circ(p-2)}\odot Bx\right)=0.$
Proof 4.2.
@@
for $j=1,2,\ldots,n$.
Notice that the introduction of $BB^{\dagger}$ in the left-hand side of (8) as follows:
$B^{\top}BB^{\dagger}\left(|Bx|^{\circ(p-2)}\odot Bx\right),$
returns the correct form at $p=2$ for a non-weighted graph, i.e.,
$B^{\top}BB^{\dagger}Bx=B^{\top}Bx=Lx.$
This motivates the choice of
@@
in which $\zeta\in\mathbb{R}_{>0}$ represents a small value. What will be referred to as the Nonlinear Modified PageRank problem can now be written in full as
@@
Remark 4.3.
The reader is strongly encouraged to review article [16] in order to help elucidate the mechanics of $f(x)$.
The Nonlinear Modified PageRank problem can reduce to the system of linear equations (6), that is, with a somewhat modified main diagonal at $p=2$. Before this can be shown, the ensuing proposition is necessary.
@@
For a connected graph with $n$ vertices, $I-B^{\dagger}B=\frac{1}{n}Q$.
Proof 4.5.
Based on [16, Theorem 1 on p. 829]. It follows from (1) that
(9) $B\left(I-B^{\dagger}B\right)=0.$
Since each row of $B$ has only two non-zero entries, those being $-1$ and $1$, the entries in each column of $\left(I-B^{\dagger}B\right)$ have to be equal for (9) to hold.
Note that $\left(I-B^{\dagger}B\right)$ fulfils the conditions of an orthogonal projection matrix, namely symmetry and idempotency. Symmetry is possible because of (4) and is the reason why all the entries of $\left(I-B^{\dagger}B\right)$ are equal. Because of the idempotent property, i.e.,
@@
the entries of $\left(I-B^{\dagger}B\right)$ are all equal to $1\big{/}n$. Otherwise, $\left(I-B^{\dagger}B\right)^{2}\neq I-B^{\dagger}B$.
@@
Let $\mathcal{F}=(\dot{\mathcal{V}},\dot{\mathcal{E}})$ be a connected and induced proper subgraph of $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$, in which $\mathcal{E}\neq\emptyset$. Assume $s\in\dot{\mathcal{V}}$ and that all adjacent vertices to $s$ are contained in $\dot{\mathcal{V}}$. Take $n=|\mathcal{V}|$, $d(v)=1\big{/}n$ with an equal contribution from the connected edges for all $v\in\mathcal{V}\setminus\dot{\mathcal{V}}$, and the number of adjacent vertices for any $v\in\mathcal{V}$ to be finite. Suppose that $\lim_{n\to\infty}\,\left(1\big{/}n\right)DQ=0$. For $p=2$, the Nonlinear Modified PageRank problem defined on $\mathcal{G}$ has the form
$T^{\infty}x=\beta r,$
where $T^{\infty}$ is a nonsingular $M$-matrix, as $n\to\infty$.
Proof 4.7.
The Nonlinear Modified PageRank problem defined on any graph $\mathcal{G}$ with $x\in\mathcal{H}(\mathcal{V})$ at $p=2$ becomes
$Tf(x)=\left(\beta D+D^{-1}L\right)f(x)=\left(\beta D+D^{-1}L\right)B^{\dagger}Bx=\left(\beta DB^{\dagger}B+D^{-1}L\right)x=\beta r.$
By Proposition 4.4,
$\left(\beta DB^{\dagger}B+D^{-1}L\right)x=\left(\beta D\left(I-\frac{1}{n}Q\right)+D^{-1}L\right)x.$
Note that $\left(1\big{/}n\right)DQ=\left(\frac{1}{\sqrt{n}}Dq\right)\left(\frac{1}{\sqrt{n}}q\right)^{\top}$. As $n\to\infty$,
$T^{\infty}x=\left(\beta D^{\infty}+D^{-1}L\right)x=\beta r.$
The superscript is to distinguish this special case. $D^{\infty}$ takes on the entries of $D$ with those on the main diagonal corresponding to $v\in\mathcal{V}\setminus\dot{\mathcal{V}}$ set equal to zero. $T^{\infty}$ is then irreducibly diagonally dominant and, therefore, nonsingular [13, Corollary 6.2.27]. It follows from [13, Corollary 6.2.27] that every eigenvalue of $T^{\infty}$ has a positive real part, owing to the real and positive entries on the main diagonal. This qualifies $T^{\infty}$ to be a nonsingular $M$-matrix [31].
Corollary 4.8.
Given $\left(T^{\infty}x\right)(v)=\beta r(v)$ for all $v\in\mathcal{V}$, a maximum value of $x$ can only occur at $s\in\mathcal{V}$ and $x(s)<1\big{/}d(s)$.
Proof 4.9.
Define
$\tilde{x}(v)=\begin{cases}x(v),&\mbox{if $v\in\mathcal{V}\setminus\dot{\mathcal{V}}$},\\ \left(\beta d(v)+1\right)x(v),&\mbox{if $v\in\dot{\mathcal{V}}\setminus\{s\}$},\\ \left(\beta d(v)+1\right)x(v)-\beta,&\mbox{if $v=s$}.\end{cases}$
It is clear from (7) that a maximum value cannot be sustained at any $v\in\mathcal{V}\setminus\dot{\mathcal{V}}$. Then the arguments in the proof of Proposition 3.3 are what is required, but will not be repeated.
In order to gain an insight into the behaviour with regard to the conductance, a perturbation argument is employed.
@@
Further to that specified in Proposition 4.6, assume $d(v)\geq 1\big{/}\left(1-\zeta\right)^{1/2}$ for all $v\in\dot{\mathcal{V}}$. For $p=2-\epsilon$, the Nonlinear Modified PageRank problem defined on $\mathcal{\mathcal{G}}$ has the form
@@
(10) $\beta r=TB^{\dagger}\left(\left((Bx)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}\odot Bx\right)\approx TB^{\dagger}\left(h\odot Bx\right),$
where $h=(1+\sigma)q+\delta$, as $n\to\infty$ and $\epsilon\to 0$. $\epsilon,\sigma\in\mathbb{R}_{\geq 0}$ along with $\delta_{i}\in\mathbb{R}$ for $i=1,2,\ldots,m$ are all small values. $\sigma$ represents an average, and $\delta$ a deviation.
Proof 4.11.
At $\epsilon=0$, the solution of problem (10) is that of the system of linear equations $T^{\infty}x=\beta r$. Note that $x$ only has non-negative entries, courtesy of $T^{\infty}$ being a nonsingular $M$-matrix and $r$ possessing only non-negative entries. Since $x(v)<1/d(s)$ for all $v\in\mathcal{V}$, entries of vector $\left((Bx)^{\circ 2}+\zeta q\right)^{\circ(1/2)}$ must be less than $1$. Consequently, as $\epsilon\to 0$, it is possible to write $h=(1+\sigma)q+\delta$.
@@
Further to that specified in Proposition 4.6, take the subgraph $\mathcal{F}$ to be non-weighted, i.e., $w(u,v)=1$ for all $\{u,v\}\in\dot{\mathcal{E}}$, and $d(v)\geq 2$ for all $v\in\dot{\mathcal{V}}$. Assume (i) there is a proper subset $\mathcal{C}$ of $\dot{\mathcal{V}}$ with $s\in\mathcal{C}$, and (ii) that $\partial(\mathcal{C})$ contains the maximum number of edges possible for any proper subset of $\dot{\mathcal{V}}$ whilst maintaining $\text{vol}\,(\mathcal{C})\leq\text{vol}\,(\mathcal{V}\setminus\mathcal{C})$. Suppose a proper subset $\mathcal{S}_{*}$ of $\dot{\mathcal{V}}$ is the sweep set based on the Nonlinear Modified PageRank problem solution at $p=2-\epsilon$, as $n\to\infty$ and $\epsilon\to 0$, which returns the smallest conductance. Proposition 4.10 provides the relevant perturbed form of the problem. The volume of subsets $\mathcal{C}$ and $\mathcal{S}_{*}$ are to be related by $\text{vol}\,(\mathcal{C})=\tau\text{vol}\,(\mathcal{S}_{*})$, in which $\tau\in\mathbb{R}_{>0}$. Let $\zeta=\left(1+\sigma\right)^{2}\tau^{2}$. If $\delta\approx 0$, $\left(\left(1\big{/}d(s)\right)^{2}+\zeta\right)^{1/2}\leq 1$ and $\text{vol}\,(\mathcal{S}_{*})\leq\text{vol}\,(\mathcal{V}\setminus\mathcal{S}_{*})$, then
$\Phi(\mathcal{S}_{*})<\left(\frac{1}{1+\sigma}\right)\Phi(\mathcal{C}).$
Remark 4.13.
$\mathcal{S}_{*}$ as a proper subset of $\dot{\mathcal{V}}$ is justifiable. Take, for example, the case where $x(v)\leq\arg\,\min_{u\in\dot{\mathcal{V}}}x(u)$ for all $v\in\mathcal{V}\setminus\dot{\mathcal{V}}$. This could potentially be accounted for in the design of the graph. Given that $d(v)=1\big{/}n$ for all $v\in\mathcal{V}\setminus\dot{\mathcal{V}}$ and $n\to\infty$, there will be no effect on the conductance if an edge formed by a vertex $u\in\mathcal{V}$ and a vertex $v\in\mathcal{V}\setminus\dot{\mathcal{V}}$ is part of the edge boundary $\partial(\mathcal{S}_{*})$.
Proof 4.14.
In the event that $\delta\approx 0$, (10) can be written as
$T^{\infty}y=\beta r,$
where $y=\left(1+\sigma\right)x$ and approximation errors are disregraded.
The cornerstone of the argument rests upon $\sum_{\{u,v\}\in\partial(\mathcal{S}_{*})}\left(\left(y(u)-y(v)\right)^{2}+\zeta\right)^{\frac{1}{2}}$. Recalling Corollary 4.8 prompts

### v2

@@ 3 PageRank problem
Further to that specified in Proposition 3.2, take $d(v)\geq 1$ for all $v\in\mathcal{V}$. Let $\mathcal{F}$ be a proper subset of $\mathcal{V}$ with $s\in\mathcal{F}$. Suppose there exists $\vartheta>0$ and $x(u)\leq\vartheta$ for all $u\in\bar{\mathcal{F}}$. If
$2\beta\vartheta\sum_{\begin{subarray}{c}u\sim v\\ u,v\in\bar{\mathcal{F}}\end{subarray}}w(u,v)\leq w\left(\partial(\mathcal{F})\right),$
then
$x[\bar{\mathcal{F}}]<\frac{1}{\beta}w\left(\partial(\mathcal{F})\right).$
Proof.
A summation of $x(v)$ over all $v\in\bar{\mathcal{F}}$ equates to
$x[\bar{\mathcal{F}}]$ $=\sum_{v\in\bar{\mathcal{F}}}\frac{1}{(\beta+1)}\sum_{u\sim v}\frac{w(u,v)}{d(u)}x(u)$
$=\frac{1}{\beta+1}\left(\sum_{\begin{subarray}{c}u\sim v\\ u,v\in\bar{\mathcal{F}}\end{subarray}}w(u,v)\left(\frac{x(u)}{d(u)}+\frac{x(v)}{d(v)}\right)+\sum_{\begin{subarray}{c}u\sim v\\ u\in\mathcal{F},v\in\bar{\mathcal{F}}\end{subarray}}\frac{w(u,v)}{d(u)}x(u)\right).$
Since $x(u)\leq\vartheta$ for all $u\in\bar{\mathcal{F}}$ and $d(v)\geq 1$ for all $v\in\mathcal{V}$, it follows that
$x[\bar{\mathcal{F}}]\leq\frac{1}{\beta+1}\left(2\vartheta\sum_{\begin{subarray}{c}u\sim v\\ u,v\in\bar{\mathcal{F}}\end{subarray}}w(u,v)+\sum_{\begin{subarray}{c}u\sim v\\ u\in\mathcal{F},v\in\bar{\mathcal{F}}\end{subarray}}w(u,v)x(u)\right).$
Based on Proposition 3.2,
$x[\bar{\mathcal{F}}]<\frac{1}{\beta+1}\left(2\vartheta\sum_{\begin{subarray}{c}u\sim v\\ u,v\in\bar{\mathcal{F}}\end{subarray}}w(u,v)+w\left(\partial(\mathcal{F})\right)\right).$
Substitution of the first term on the right-hand side with $w\left(\partial(\mathcal{F})\right)/\beta$ leads to
$x[\bar{\mathcal{F}}]<\frac{1}{\beta+1}\left(\frac{1}{\beta}w\left(\partial(\mathcal{F})\right)+w\left(\partial(\mathcal{F})\right)\right)=\frac{1}{\beta}w\left(\partial(\mathcal{F})\right).\qed$
This demonstrates that for any proper subset $\mathcal{F}$ containing $s$ and satisfying the stated condition, the cut weight scaled by $1/\beta$ bounds the sum value of PageRank vector entries corresponding to vertices contained in the complement $\bar{\mathcal{F}}$.
4 Nonlinear PageRank problem
@@
is proposed for local graph partitioning purposes. The question is what form should $f$ take.
@@
Let $p\in(1,\infty)$. Then
$\frac{\partial}{\partial x}\left(\lVert Bx\rVert_{p}^{p}\right)=p\left(|Bx|^{\circ(p-2)}\odot Bx\right)^{\top}B.$
Proof.
@@
for $j=1,2,\ldots,n$, where summation over the repeated index is assumed. ∎
If $\frac{\partial}{\partial x}\left(\lVert Bx\rVert_{2}^{2}\right)=0$, then $B^{\top}Bx=0$. Observe that, by (1a),
$B^{\top}Bx=Lx=B^{\top}BB^{\dagger}Bx=LB^{\dagger}Bx$
for a non-weighted graph. This inspires the choice of
@@
in which $\zeta\in\mathbb{R}_{>0}$ represents a small value. What will be referred to as the Nonlinear PageRank problem can now be written in full as
@@
Remark 2.
Properties of the Moore-Penrose inverse of an arbitrary incidence matrix are examined in [25].
The solution to the Nonlinear PageRank problem for $p=2$ can be expressed as an additive combination of the corresponding (linear) PageRank problem solution and a constant term. Before this can be demonstrated, the following proposition must first be established.
@@
For a simple and connected graph with $n$ vertices, $I-B^{\dagger}B=\frac{1}{n}Q$.
Proof.
Based on [25, Theorem 1 on p. 829]. It follows from (1a) that
$B\left(I-B^{\dagger}B\right)=0.$ (4)
Since each row of $B$ has only two non-zero entries, those being 1-1 and $1$, the entries in each column of $\left(I-B^{\dagger}B\right)$ have to be equal for (4) to hold.
Note that $\left(I-B^{\dagger}B\right)$ fulfils the conditions of an orthogonal projection matrix, namely symmetry and idempotency. Symmetry is possible because of (1d) and is the reason why all the entries of $\left(I-B^{\dagger}B\right)$ are equal. Because of the idempotent property, i.e.,
@@
the entries of $\left(I-B^{\dagger}B\right)$ are all equal to $1\big/n$. Otherwise, $\left(I-B^{\dagger}B\right)^{2}\neq I-B^{\dagger}B$. ∎
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ with $\mathcal{E}\neq\emptyset$ and $n=|\mathcal{V}|$. For $p=2$, the solution in the least squares sense to the Nonlinear PageRank problem defined on $\mathcal{G}$ has the form
$x=c-\frac{1}{n}q,$
where $c=\beta T^{-1}r$ is the solution of the corresponding (linear) PageRank problem.
Proof.
The Nonlinear PageRank problem for $p=2$ defined on $\mathcal{G}$ with $x\in\mathcal{H}(\mathcal{V})$ equates to
$TB^{\dagger}Bx=\beta r.$
Proposition 3.1 discloses the nonsingular nature of $T$, which supports $B^{\dagger}Bx=c$. By Proposition 2.3, the minimal least squares solution then is
$x=B^{\dagger}Bc.$
This can be re-expressed as
$x=\left(I-\frac{1}{n}Q\right)c=c-\frac{1}{n}qq^{\top}c$
courtesy of Proposition 4.2. Recall that $\alpha=1/(1+\beta)$ and $\beta\in(0,1)$. As such, $q^{\top}c=1$. Thus, x=c(1/n)qx=c-(1/n)q. ∎
Remark 3.
Clearly, $\lim_{n\to\infty}\lVert(1/n)q\rVert=\lim_{n\to\infty}1/\sqrt{n}=0$.
In order to gain an insight into the effect that the Nonlinear PageRank problem or rather the solution has on the conductance, a perturbation argument is developed.
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ with $\mathcal{E}\neq\emptyset$ and $n=|\mathcal{V}|$. Consider the case where $0<w_{\min}\leq w(v_{i},v_{j})\leq w_{\max}<\infty$ for all $\{v_{i},v_{j}\}\in\mathcal{E}$ and every vertex $v\in\mathcal{V}$ has at most $\Delta$ neighbours. $x^{0}$ is to be the least squares solution of the Nonlinear PageRank problem defined on $\mathcal{G}$ at $p=2$. Take $\epsilon\in\mathbb{R}_{\geq 0}$ to be a small value. For sufficiently large (but finite) $n$, there exists $x(\epsilon)=x^{0}+\epsilon x^{1}+\mathcal{O}(\epsilon^{2})$ such that the Perturbed Nonlinear PageRank problem
@@
$TB^{\dagger}\left(\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}\odot Bx(\epsilon)\right)=\beta r$ (5)
satisfies
$\left\lVert\beta r-TB^{\dagger}\left(\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}\odot Bx(\epsilon)\right)\right\rVert\leq\frac{h_{1}}{\sqrt{n}}+\frac{h_{2}}{\sqrt{n}}\epsilon+h_{3}\epsilon^{2},$
in which $h_{1}=(\beta+1)+\sqrt{(w_{max}/w_{min})\Delta}$ and $h_{2},h_{3}>0$ are constants independent of both $\epsilon$ and $n$.
Proof.
A Taylor series expansion of $\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}$ truncated at the second order yields
$\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}=q-\frac{\epsilon}{2}y+\mathcal{O}(\epsilon^{2}),$
where $y=\ln\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)$.
Since $x(\epsilon)=x^{0}+\epsilon x^{1}+\mathcal{O}(\epsilon^{2})$, it follows that
$\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}\odot Bx(\epsilon)$ $=\left(q-\frac{\epsilon}{2}y+\mathcal{O}(\epsilon^{2})\right)\odot\left(Bx^{0}+\epsilon Bx^{1}+\mathcal{O}(\epsilon^{2})\right)$
$=Bx^{0}+\epsilon Bx^{1}-\frac{\epsilon}{2}\left(y\odot Bx^{0}\right)+\mathcal{O}(\epsilon^{2})$
and
$TB^{\dagger}\left(\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}\odot Bx(\epsilon)\right)=TB^{\dagger}Bx^{0}+\epsilon TB^{\dagger}Bx^{1}\\ -\frac{\epsilon}{2}TB^{\dagger}\left(y\odot Bx^{0}\right)+\mathcal{O}(\epsilon^{2}).$
By Propositions 4.2 and 4.3,
$TB^{\dagger}Bx^{0}$ $=T\left(I-\frac{1}{n}qq^{\top}\right)\left(c-\frac{1}{n}q\right)$
$=T\left(c-\frac{1}{n}qq^{\top}c-\frac{1}{n}q+\frac{1}{n^{2}}qq^{\top}q\right)$
$=Tc-\frac{1}{n}Tq,$
as $q^{\top}c=1$. But $c=\beta T^{-1}r$, which leads to $TB^{\dagger}Bx^{0}=\beta r-(1/n)Tq$.
It is now possible to re-express the Perturbed Nonlinear PageRank problem in the form
$\beta r=\beta r-\frac{1}{n}Tq+\epsilon TB^{\dagger}Bx^{1}-\frac{\epsilon}{2}TB^{\dagger}\left(y\odot Bx^{0}\right)+\mathcal{O}(\epsilon^{2}).$
Set $x^{1}=(1/2)B^{\dagger}\left(y\odot Bx^{0}\right)$ so that $TB^{\dagger}Bx^{1}=(1/2)TB^{\dagger}\left(y\odot Bx^{0}\right)$. Then
$\left\lVert\beta r-TB^{\dagger}\left(\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}\odot Bx(\epsilon)\right)\right\rVert\leq\frac{1}{n}\left\lVert Tq\right\rVert+\left\lVert\mathcal{O}(\epsilon^{2})\right\rVert.$
Recall that $T=\beta I+LD^{-1}=\beta I+(D-A)D^{-1}=(\beta+1)I-AD^{-1}$. This leads to
$\left\lVert Tq\right\rVert\leq\left\lVert T\right\rVert\left\lVert q\right\rVert=\left\lVert(\beta+1)I-AD^{-1}\right\rVert\sqrt{n}\leq(\beta+1)\sqrt{n}+\left\lVert AD^{-1}\right\rVert\sqrt{n}.$
Based on $\left\lVert AD^{-1}\right\rVert\leq\left\lVert AD^{-1}\right\rVert_{1}^{\frac{1}{2}}\left\lVert AD^{-1}\right\rVert_{\infty}^{\frac{1}{2}}=\sqrt{(w_{max}/w_{min})\Delta}$,
$\frac{1}{n}\left\lVert Tq\right\rVert\leq\frac{1}{\sqrt{n}}(\beta+1)+\frac{1}{\sqrt{n}}\left(\frac{w_{max}}{w_{min}}\Delta\right)^{\frac{1}{2}}=\frac{1}{\sqrt{n}}\left((\beta+1)+\left(\frac{w_{max}}{w_{min}}\Delta\right)^{\frac{1}{2}}\right).$
Finally,
$\left\lVert\beta r-TB^{\dagger}\left(\left(\left(Bx^{0}\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}\odot Bx(\epsilon)\right)\right\rVert\leq\frac{h_{1}}{\sqrt{n}}+\frac{h_{2}}{\sqrt{n}}\epsilon+h_{3}\epsilon^{2}.\qed$
Notice that (5) does not exactly match the Nonlinear PageRank problem at $p=2-\epsilon$ with $x=x(\epsilon)$, which has the form
$TB^{\dagger}\left(\left(\left(Bx(\epsilon)\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}\odot Bx(\epsilon)\right)=\beta r.$
There is $Bx^{0}$ in place of a $Bx(\epsilon)$ term. However, this does not detract from the value of the Perturbed Nonlinear PageRank problem as a representation of the Nonlinear PageRank problem at $p=2-\epsilon$ with $x=x(\epsilon)$. $((Bx^{0})^{\circ 2}+\zeta q)^{\circ-\frac{\epsilon}{2}}$ as a stand-in for
$\left(\left(Bx(\epsilon)\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}$ $=\left(\left(Bx^{0}+\epsilon Bx^{1}+\mathcal{O}(\epsilon^{2})\right)^{\circ 2}+\zeta q\right)^{\circ-\frac{\epsilon}{2}}$
$=\left(\left(Bx^{0}\right)^{\circ 2}+2\left(Bx^{0}\odot\epsilon Bx^{1}\right)+\mathcal{O}(\epsilon^{2})+\zeta q\right)^{\circ-\frac{\epsilon}{2}}$
is a justified simplification when $\epsilon\ll 1$.
The expansion $x(\epsilon)=x^{0}+\epsilon x^{1}+\mathcal{O}(\epsilon^{2})$ provides a means by which to gain an appreciation as to the impact of $p$ in the Nonlinear PageRank problem. This is captured in the following proposition. The conductance obtained from the solution of the Perturbed Nonlinear PageRank problem is compared to that of a reference set $\mathcal{F}$.
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},w)$ with $\mathcal{E}\neq\emptyset$ and $n=|\mathcal{V}|$. Take $w(u,v)=1$ for all $\{u,v\}\in\mathcal{E}$. Every vertex $v\in\mathcal{V}$ has at most $\Delta$ neighbours. $\mathcal{F}$ is to be a proper subset of $\mathcal{V}$ containing the starting vertex $s$ such that
  • $|\partial(\mathcal{F})|$ is a maximum and
  • no edge of $\partial(\mathcal{F})$ includes $s$.
$x^{0}$ is to be the least squares solution of the Nonlinear PageRank problem defined on $\mathcal{G}$ at $p=2$. Assume that $(Bx^{0})(v)+\zeta\leq 1$ for all $v\in\mathcal{V}$ and $0\leq\epsilon\ll 1$. Consider the case where $n>h_{1}^{2}/\epsilon^{4}$ and $((Bx^{0})^{\circ 2}+\zeta q)^{\circ-\frac{\epsilon}{2}}$ admits the decomposition